### v3

# A Probabilistic Model for Forest Fires

Abstract
We propose a discrete two-dimensional mathematical model for forest fires and we derive certain results describing its limiting behavior. We also pose a relevant open question.
Keywords. Random field; limiting behavior.
2010 AMS Mathematics Classification. 60G60; 92F05.
1 Introduction
The forest is represented by the set
$\mathbb{N}^{2}=\{(j,k)\,:\,j,k\in\mathbb{N}\},$
where $\mathbb{N}=\{0,1,2,\ldots\}$ is the set of nonnegative integers. Each point $(j,k)\in\mathbb{N}^{2}$ is associated to a tree. The status of the tree at $(j,k)$ is denoted by $X(j,k)$ (later we will use additional notations for $X(j,k)$), and it takes two values, 0 or $1$, so that $X(j,k)=1$ means that the tree at $(j,k)$ is burnt, while $X(j,k)=0$ means that the tree at $(j,k)$ is not burnt.
The key feature of the model is the following: The status of the tree at $(j,k)$ is affected only by the status of the neighboring trees at (j1,k)(j-1,k) and (j,k1)(j,k-1) so that
(i) if the trees at (j1,k)(j-1,k) and (j,k1)(j,k-1) are not burnt, then tree at $(j,k)$ is not burnt;
(ii) if the tree at (j1,k)(j-1,k) is burnt and the tree at (j,k1)(j,k-1) is not burnt, then there is a probability $\alpha$ that the tree at $(j,k)$ is burnt;
(iii) if the tree at (j1,k)(j-1,k) is not burnt and the tree at (j,k1)(j,k-1) is burnt, then there is a probability $\beta$ that the tree at $(j,k)$ is burnt;
(iv) if both trees at (j1,k)(j-1,k) and (j,k1)(j,k-1) are burnt, then there is a probability $\gamma$ that the tree at $(j,k)$ is burnt.
To avoid paradoxical situations we need to assume that $\gamma\geq\max\{\alpha,\beta\}$, since one expects that in the case where both trees at (j1,k)(j-1,k) and (j,k1)(j,k-1) are burnt there are more chances that the tree at $(j,k)$ is burnt than in the case where only one of the trees at (j1,k)(j-1,k) and (j,k1)(j,k-1) is burnt.
The probabilities $\alpha,\beta,\gamma$ are related to the wind speed and wind direction.
We assume that the fire starts at the point $(0,0)$ so that $X(0,0)=1$.
The complete description of the model is given in the next section.
The main results of the paper concern the sequence of random variables
$Y_{n}=X(0,n)+X(1,n-1)+\cdots+X(n,0)$
since this sequence contains significant information regarding the fate of the forest.
In the case $\alpha+\beta<1$ and $\gamma<1$ we show that $\sum_{n=0}^{\infty}Y_{n}<\infty$ almost surely, which implies that the forest is not considerably damaged.
The case $\alpha+\beta=\gamma=1$, with $\alpha\beta>0$, is a borderline case. Here we show that $Y_{n}\to 0$ almost surely, but $\mathbb{E}[Y_{n}]=1$ for all $n\geq 0$. This gives a somehow “mixed” information regarding the fate of the forest.
The case $\alpha+\beta>1$ is the most challenging. Here, with the additional hypothesis that $\gamma=1$ we show that
$\mathbb{P}\left\{Y_{n}=0\right\}\leq\frac{n}{n+1}(2-\alpha-\beta)\qquad\text{and}\qquad\mathbb{E}\left[Y_{n}\right]\geq 1+(\alpha+\beta-1)n,\qquad n\geq 0.$
These estimates indicate that the fire is quite catastrophic.
In the opposite direction, we have the bounds
$\lim_{n}\mathbb{P}\left\{Y_{n}=0\right\}\geq\frac{(1-\alpha)(1-\beta)}{\alpha\beta}\qquad\text{and}\qquad\limsup_{n}\frac{\mathbb{E}\left[Y_{n}\right]}{n}\leq\frac{\alpha+\beta-1}{\alpha\beta},$
which are, clearly, also valid in the case where $\gamma<1$ (as long as $\alpha+\beta>1$).
Finally, in the case $\alpha+\beta>1$ an interesting open question is whether $Y_{n}/n$ converges, at least in distribution.
2 The law of the random field $X(j,k)$
In the sequel we will use the notations
$X_{j}^{j+k}=X_{jk}=X(j,k),\qquad\text{so that}\qquad X_{j}^{n}=X(j,n-j),$ (1)
e.g., $X_{2}^{5}=X_{23}=X(2,3)$. Thus, for convenience, sometimes $X(j,k)$ will be denoted by $X_{jk}$ and sometimes by $X_{j}^{n}$, where $n=j+k$.
The “boundary conditions” for $X(j,k)$ are
$X(0,0)=1\qquad\text{and}\qquad X(j,-1)=X(-1,k)=0\qquad\text{for all }j,k\in\mathbb{N}$ (2)
(we can further assume that $X(j,k)=0$ whenever $j<0$ or $k<0$).
We will now define the random field $X(j,k)$, $j,k\in\mathbb{N}$, inductively.
Notice that $X(j,k)$ is a Bernoulli random variable, i.e. it takes only the values 0 and $1$.
Start with the boundary conditions (2). Next, suppose that for some $n\geq 0$ we have defined the field $X_{j}^{j+k}=X_{jk}$ for $0\leq j+k\leq n$. We will show how to define (the law of) $X_{j}^{j+k}$ for $j+k=n+1$.
First we introduce the $\sigma$-algebras
$\mathcal{F}_{n}=\sigma\left(X_{jk},\ 0\leq j+k\leq n\right),\qquad\mathcal{G}_{n}=\sigma\left(X_{jk},\ j+k=n\right),\qquad n\geq 0$ (3)
(of course, in view of (2), we have that $\mathcal{F}_{0}=\mathcal{G}_{0}=\{\varnothing,\Omega\}$).
For each (fixed) $j\in\{0,1,\ldots,n+1\}$ we define the law of $X_{j}^{n+1}$ by setting
$\left.\begin{array}[]{l}\mathbb{P}\big\{X_{j}^{n+1}=1\,\big|\,X_{j-1}^{n}=X_{j}^{n}=0\big\}=0,\\ \mathbb{P}\big\{X_{j}^{n+1}=1\,\big|\,X_{j-1}^{n}=1,\ X_{j}^{n}=0\big\}=\alpha,\\ \mathbb{P}\big\{X_{j}^{n+1}=1\,\big|\,X_{j-1}^{n}=0,\ X_{j}^{n}=1\big\}=\beta,\\ \mathbb{P}\big\{X_{j}^{n+1}=1\,\big|\,X_{j-1}^{n}=X_{j}^{n}=1\big\}=\gamma,\end{array}\right\}$ (4)
where
$0\leq\alpha,\beta\leq\gamma\leq 1$ (5)
($\alpha$, $\beta$, and $\gamma$ are given). The set of equations in (4) presents the main feature of the model, namely that the status of the tree at $(j,k)$ is affected only by the status of the neighboring trees at (j1,k)(j-1,k) and (j,k1)(j,k-1).
Now, the joint distribution of $X_{j}^{n+1}$, $0\leq j\leq n+1$, is defined by setting
$\mathbb{P}\big\{X_{j}^{n+1}=\epsilon_{j},\ 0\leq j\leq n+1\,\big|\,\mathcal{F}_{n}\big\}$ $=\mathbb{P}\big\{X_{j}^{n+1}=\epsilon_{j},\ 0\leq j\leq n+1\,\big|\,\mathcal{G}_{n}\big\}$
$=\prod_{j=0}^{n+1}\mathbb{P}\big\{X_{j}^{n+1}=\epsilon_{j}\,\big|\,\mathcal{G}_{n}\big\}$
$=\prod_{j=0}^{n+1}\mathbb{P}\big\{X_{j}^{n+1}=\epsilon_{j}\,\big|\,X_{j-1}^{n},X_{j}^{n}\big\},$ (6)
where each $\epsilon_{j}$, $0\leq j\leq n+1$, can be either 0 or $1$ (recall that X(1,n)X(-1,n) and X(n,1)X(n,-1) are deterministic, i.e. they have given nonrandom values).
The second equality in (6) states that the random variables $X_{j}^{n+1}$, $0\leq j\leq n+1$, are conditionally independent given $\mathcal{G}_{n}$. Actually, since (6) implies that
$\mathbb{P}\big\{X_{j}^{n+1}=\epsilon_{j}\,\big|\,\mathcal{G}_{n}\big\}=\mathbb{P}\big\{X_{j}^{n+1}=\epsilon_{j}\,\big|\,\mathcal{F}_{n}\big\},$
it follows that $X_{j}^{n+1}$, $0\leq j\leq n+1$, are also conditionally independent given $\mathcal{F}_{n}$ (for instance, since $\mathcal{F}_{0}$ is the trivial $\sigma$-algebra, the random variables $X_{0}^{1}$ and $X_{1}^{1}$ are independent).
Let us also notice that the defining equations in (6) have a Markovian flavor [2].
One consequence of the definition of the random field $X_{j}^{n}$ is that there is a subset $\Omega_{0}$ of $\Omega$ with $\mathbb{P}(\Omega_{0})=1$ such that
$\left\{X_{j}^{n}=0,\ 0\leq j\leq n\right\}\cap\Omega_{0}\subset\left\{X_{j}^{n+1}=0,\ 0\leq j\leq n+1\right\}\cap\Omega_{0},\qquad n\geq 0.$ (7)
Remark 1. From the definition of the random field $X(j,k)$ it follows that if
$\tilde{\alpha}\geq\alpha,\qquad\tilde{\beta}\geq\beta,\qquad\tilde{\gamma}\geq\gamma$
and $\tilde{X}(j,k)$ is the random field associated to $\tilde{\alpha},\tilde{\beta},\tilde{\gamma}$, while $X(j,k)$ is the random field associated to $\alpha,\beta,\gamma$, then
$\mathbb{P}\left\{\tilde{X}(j,k)=1\right\}\geq\mathbb{P}\left\{X(j,k)=1\right\}\qquad\text{for every }j,k\geq 0,$ (8)
i.e. $\tilde{X}(j,k)$ is stochastically larger than $X(j,k)$ for every $j,k\geq 0$. $\diamondsuit$
The symbol $\diamondsuit$ indicates the end of a remark or an example.
Remark 2. From formulas (2) and (4) it follows that if $\alpha=\beta=1$, then $X(j,k)=1$ a.s. for all $(j,k)\in\mathbb{N}^{2}$, i.e. all trees of the forest are burnt, while if $\alpha=\beta=0$, then $X(j,k)=0$ a.s. for all $(j,k)\in\mathbb{N}^{2}\smallsetminus\{(0,0)\}$, i.e. the only burnt tree is the one at $(0,0)$.
If $\alpha=0$, then (2) and (4) imply that the only trees that are possibly burnt are located at the points $(0,k)$, $k\in\mathbb{N}$. Likewise, if $\beta=0$, then (2) and (4) imply that the only trees that are possibly burnt are located at the points $(j,0)$, $j\in\mathbb{N}$. Therefore, if $\alpha\beta=0$, our model is one-dimensional, and, consequently, very simple. We analyze this model in the Appendix. $\diamondsuit$
For the rest of the paper we will always assume that
$\alpha>0\qquad\text{and}\qquad\beta>0.$ (9)
Next, we present some immediate consequences of the defining formulas (4) and (6). But, first, let us recall that for a Bernoulli random variable $X$ we have that
${\bf 1}_{\{X=1\}}=X,\qquad{\bf 1}_{\{X=0\}}=1-X,\quad\text{and}\quad\mathbb{E}\left[z^{X}\right]=1+\mathbb{E}[X](z-1).$ (10)
From (4), (6), and (10) it follows that
$\mathbb{P}\left\{X_{j}^{n}=1\,\big|\,\mathcal{F}_{n-1}\right\}$ $=\mathbb{E}\left[X_{j}^{n}\,\big|\,\mathcal{F}_{n-1}\right]$
$=\mathbb{E}\left[X_{j}^{n}\,\big|\,X_{j-1}^{n-1},X_{j}^{n-1}\right]$
$=\alpha X_{j-1}^{n-1}\left(1-X_{j}^{n-1}\right)+\beta X_{j}^{n-1}\left(1-X_{j-1}^{n-1}\right)+\gamma X_{j-1}^{n-1}X_{j}^{n-1}$
$=\alpha X_{j-1}^{n-1}+\beta X_{j}^{n-1}-(\alpha+\beta-\gamma)X_{j-1}^{n-1}X_{j}^{n-1}.$ (11)
Also, from (10) and (11) we get
$\mathbb{E}\left[z^{X_{j}^{n}}\,\big|\,\mathcal{F}_{n-1}\right]$ $=1+\mathbb{E}\left[X_{j}^{n}\,\big|\,\mathcal{F}_{n-1}\right](z-1)$
$=1+\left[\alpha X_{j-1}^{n-1}+\beta X_{j}^{n-1}-(\alpha+\beta-\gamma)X_{j-1}^{n-1}X_{j}^{n-1}\right](z-1).$ (12)
Finally, from (11) and the fact that, for $j\neq k$, the random variables $X_{j}^{n}$ and $X_{k}^{n}$ are conditionally independent given $\mathcal{F}_{n-1}$ we get
$\mathbb{E}\left[X_{j}^{n}X_{k}^{n}\,\big|\,\mathcal{F}_{n-1}\right]$ $=\mathbb{E}\left[X_{j}^{n}\,\big|\,\mathcal{F}_{n-1}\right]\mathbb{E}\left[X_{k}^{n}\,\big|\,\mathcal{F}_{n-1}\right]$
$=\left[\,\alpha X_{j-1}^{n-1}+\beta X_{j}^{n-1}-(\alpha+\beta-\gamma)X_{j-1}^{n-1}X_{j}^{n-1}\right]$
$\quad\times\left[\,\alpha X_{k-1}^{n-1}+\beta X_{k}^{n-1}-(\alpha+\beta-\gamma)X_{k-1}^{n-1}X_{k}^{n-1}\right].$ (13)
At the end of this section we give few examples which indicate how to calculate certain probabilities related to the field $X(j,k)$.
Example 1. For a given $n\geq 0$ let us show how to compute the probability
$\mathbb{P}\left\{X_{j}^{n}=\epsilon_{j},\ 0\leq j\leq n\right\},$ (14)
where the value of each $\epsilon_{j}$, $0\leq j\leq n$, is given (it is either 0 or $1$).
For $n=0$ we have that $X_{0}^{0}=1$ is deterministic.
For $n=1$ the probability in (14) becomes, in view of (6) and the fact that $\mathcal{G}_{0}$ is the trivial $\sigma$-algebra
$\mathbb{P}\left\{X_{0}^{1}=\epsilon_{0},\ X_{1}^{1}=\epsilon_{1}\right\}=\mathbb{P}\left\{X_{0}^{1}=\epsilon_{0}\right\}\mathbb{P}\left\{X_{1}^{1}=\epsilon_{1}\right\}$ (15)
and these probabilities can be computed from the boundary conditions together with the equations of (4).
Suppose that for $n\leq m$ we can compute the probability in (14) by using (6) and (4) (and the boundary condition). We will, then, show how to calculate the probability in (14) for $n=m+1$. We have
$\mathbb{P}\left\{X_{j}^{m+1}=\epsilon_{j},\ 0\leq j\leq m+1\right\}$
$=\sum_{\tilde{\epsilon}_{k}=0\text{ or }1\atop 0\leq k\leq m}\left[\,\mathbb{P}\left\{X_{j}^{m+1}=\epsilon_{j},\ 0\leq j\leq m+1\,\big|\,X_{k}^{m}=\tilde{\epsilon}_{k},\ 0\leq k\leq m\right\}\right$.
$\qquad\qquad\ \ \left.\times\,\mathbb{P}\left\{X_{k}^{m}=\tilde{\epsilon}_{k},\ 0\leq k\leq m\right\}\right],$ (16)
where the sum is taken over all the $2^{m+1}$ different choices of $\tilde{\epsilon}_{k}$, $0\leq k\leq m$.
Each term of the sum in (16) is a product of a conditional probability and a probability. The latter can be computed by the induction hypothesis. As for the conditional probability, it can be computed too by using (6) (actually, in view of (6), the events $\{X_{j}^{m+1}=\epsilon_{j}\}$, $0\leq j\leq m+1$, are conditionally independent given the event $\{X_{k}^{m}=\tilde{\epsilon}_{k},\ 0\leq k\leq m\}$). This finishes the induction which shows how all probabilities of the form of (14) can be computed. $\diamondsuit$
Example 2. Let us compute the probability
$\mathbb{P}\big\{X_{12}=0,\ X_{11}=1\big\}.$ (17)
We have
$\mathbb{P}\big\{X_{12}=0,\ X_{11}=1\big\}=\sum_{\epsilon=0,1}\mathbb{P}\big\{X_{12}=0\,\big|\,X_{02}=\epsilon,\ X_{11}=1\big\}\mathbb{P}\big\{X_{02}=\epsilon,\ X_{11}=1\big\},$
where all quantities in the right-hand side can be computed. Indeed, (4) yields
$\mathbb{P}\big\{X_{12}=0\,\big|\,X_{02}=0,\ X_{11}=1\big\}=1-\beta$
and
$\mathbb{P}\big\{X_{12}=0\,\big|\,X_{02}=1,\ X_{11}=1\big\}=1-\gamma,$
while the probability $\mathbb{P}\{X_{02}=\epsilon,\ X_{11}=1\}$ can be computed with the help of Example 1. $\diamondsuit$
Example 3. Let us also compute the probability
$\mathbb{P}\big\{X_{22}=1,\ X_{12}=0,\ X_{11}=1\big\}.$ (18)
We have
$\mathbb{P}\big\{X_{22}=1,\ X_{12}=0,\ X_{11}=1\big\}$
$=$ $\sum_{\epsilon=0,1}\mathbb{P}\big\{X_{22}=1\,\big|\,X_{12}=0,\ X_{21}=\epsilon,\ X_{11}=1\big\}\mathbb{P}\big\{X_{12}=0,\ X_{21}=\epsilon,\ X_{11}=1\big\},$
$=$ $\sum_{\epsilon=0,1}\mathbb{P}\big\{X_{22}=1\,\big|\,X_{12}=0,\ X_{21}=\epsilon\big\}\mathbb{P}\big\{X_{12}=0,\ X_{21}=\epsilon,\ X_{11}=1\big\}$
$=$ $\,\mathbb{P}\big\{X_{22}=1\,\big|\,X_{12}=0,\ X_{21}=1\big\}\mathbb{P}\big\{X_{12}=0,\ X_{21}=1,\ X_{11}=1\big\}$
$=$ $\,\beta\,\mathbb{P}\big\{X_{12}=0,\ X_{21}=1,\ X_{11}=1\big\},$
where the second equality is a consequence of (6), while the last two equalities follow from (4).
Finally,
$\mathbb{P}\big\{X_{12}=0,\ X_{21}=1,\ X_{11}=1\big\}$
$=\sum_{\epsilon_{0},\epsilon_{2}=0,1}\left[\,\mathbb{P}\big\{X_{12}=0,\ X_{21}=1\,\big|\,X_{02}=\epsilon_{0},\ X_{11}=1,\ X_{20}=\epsilon_{2}\big\}\right$.
$\qquad\qquad\ \left.\times\,\mathbb{P}\big\{X_{02}=\epsilon_{0},\ X_{11}=1,\ X_{20}=\epsilon_{2}\big\}\right]$
$=\sum_{\epsilon_{0},\epsilon_{2}=0,1}\left[\,\mathbb{P}\big\{X_{12}=0\,\big|\,X_{02}=\epsilon_{0},\ X_{11}=1\big\}\mathbb{P}\big\{X_{21}=1\,\big|\,X_{11}=1,\ X_{20}=\epsilon_{2}\big\}\right$.
$\qquad\qquad\ \left.\times\,\mathbb{P}\big\{X_{02}=\epsilon_{0},\ X_{11}=1,X_{20}=\epsilon_{2}\big\}\right]$
(the last equality follows from (6), namely the conditional independence given $\mathcal{G}_{2}$), where all probabilities can be computed with the help of (4) and Example 1. $\diamondsuit$
Example 4. Here we compute the conditional probabilities
$w_{k}=\mathbb{P}\left\{X_{1}^{k}=0\,\big|\,X_{0}^{1}=X_{0}^{2}=\cdots=X_{0}^{k}=1\right\},\qquad k\geq 1.$ (19)
For $k=1$ we have (since $X_{0}^{1}$ and $X_{1}^{1}$ are independent)
$w_{1}=\mathbb{P}\left\{X_{1}^{1}=0\,\big|\,X_{0}^{1}=1\right\}=\mathbb{P}\left\{X_{1}^{1}=0\right\}=1-\alpha.$ (20)
For $k\geq 2$ the conditional probability of (19) can be expressed as
$w_{k}=\frac{\mathbb{P}\left\{X_{0}^{k}=1,X_{1}^{k}=0\,\big|\,X_{0}^{1}=X_{0}^{2}=\cdots=X_{0}^{k-1}=1\right\}}{\mathbb{P}\left\{X_{0}^{k}=1\,\big|\,X_{0}^{1}=X_{0}^{2}=\cdots=X_{0}^{k-1}=1\right\}},$ (21)
where from the defining properties of the random field $X_{j}^{n}$ we have that
$\mathbb{P}\left\{X_{0}^{k}=1\,\big|\,X_{0}^{1}=X_{0}^{2}=\cdots=X_{0}^{k-1}=1\right\}=\mathbb{P}\left\{X_{0}^{k}=1\,\big|\,X_{0}^{k-1}=1\right\}=\beta.$ (22)
hence (21) becomes
$w_{k}=\beta^{-1}\mathbb{P}\left\{X_{0}^{k}=1,X_{1}^{k}=0\,\big|\,A_{k-1}\right\},$ (23)
where for typographical convenience we have set
$A_{k}=\left\{X_{0}^{1}=X_{0}^{2}=\cdots=X_{0}^{k}=1\right\},\qquad k\geq 1.$ (24)
Now,
$\mathbb{P}\left\{X_{0}^{k}=1,X_{1}^{k}=0\,\big|\,A_{k-1}\right\}$
$=\mathbb{P}\left\{X_{0}^{k}=1,X_{1}^{k}=0\,\big|\,A_{k-1},X_{1}^{k-1}=0\right\}\mathbb{P}\left\{X_{1}^{k-1}=0\,\big|\,A_{k-1}\right\}$
$\quad+\mathbb{P}\left\{X_{0}^{k}=1,X_{1}^{k}=0\,\big|\,A_{k-1},X_{1}^{k-1}=1\right\}\mathbb{P}\left\{X_{1}^{k-1}=1\,\big|\,A_{k-1}\right\}.$ (25)
In view of (6), (19), and (24), formula (25) becomes
$\mathbb{P}\left\{X_{0}^{k}=1,X_{1}^{k}=0\,\big|\,A_{k-1}\right\}$
$=\mathbb{P}\left\{X_{0}^{k}=1,X_{1}^{k}=0\,\big|\,X_{0}^{k-1}=1,X_{1}^{k-1}=0\right\}w_{k-1}$
$\quad+\mathbb{P}\left\{X_{0}^{k}=1,X_{1}^{k}=0\,\big|\,X_{0}^{k-1}=1,X_{1}^{k-1}=1\right\}(1-w_{k-1})$
$=\mathbb{P}\left\{X_{0}^{k}=1\,\big|\,X_{0}^{k-1}=1\right\}\mathbb{P}\left\{X_{1}^{k}=0\,\big|\,X_{0}^{k-1}=1,X_{1}^{k-1}=0\right\}w_{k-1}$
$\quad+\mathbb{P}\left\{X_{0}^{k}=1\,\big|\,X_{0}^{k-1}=1\right\}\mathbb{P}\left\{X_{1}^{k}=0\,\big|\,X_{0}^{k-1}=1,X_{1}^{k-1}=1\right\}(1-w_{k-1}).$ (26)
Thus, by using (4) in (26) we obtain
$\mathbb{P}\left\{X_{0}^{k}=1,X_{1}^{k}=0\,\big|\,A_{k-1}\right\}=\beta(1-\alpha)w_{k-1}+\beta(1-\gamma)(1-w_{k-1}),$ (27)
and, finally, in view of (27), formula (23) yields
$w_{k}=(1-\alpha)w_{k-1}+(1-\gamma)(1-w_{k-1}).$ (28)
Formula (28) is a linear first-order difference equation with constant coefficients, whose initial condition is (20). Therefore, it is easy to get that
$w_{k}=\frac{\alpha}{1-(\gamma-\alpha)}(\gamma-\alpha)^{k}+\frac{1-\gamma}{1-(\gamma-\alpha)},\qquad k\geq 1.$ (29)
$\diamondsuit$
2.1 An equivalent Markovian model
The above forest fire model can be viewed as a Markov chain.
Let us start with a definition.
Definition 1. A binary sequence is a sequence $b=\{b_{j}\}_{j=0}^{\infty}$ such that $b_{j}=0$ or $1$ for every $j\geq 0$.
The valence $\boldsymbol{v}(b)$ of a binary sequence $b=\{b_{j}\}_{j=0}^{\infty}$ is the number of its elements which are equal to $1$, namely
$\boldsymbol{v}(b)=\sum_{j=0}^{\infty}b_{j}.$ (30)
The set of all binary sequences of finite valence is denoted by $\mathcal{S}$. Thus
$\mathcal{S}=\left\{b\,:\,b\text{ is a binary sequence with }\boldsymbol{v}(b)<\infty\right\}.$ (31)
Next, we define a Markov chain $\hat{X}(n)$, $n\in\mathbb{N}$, whose state space is the set $\mathcal{S}$ of (31). Thus, for each $n\in\mathbb{N}$ we have that
$\hat{X}(n)=\left\{\hat{X}_{j}(n)\right\}_{j=0}^{\infty}\quad\text{with}\quad\hat{X}_{j}(n)=0\text{ or }1\quad\text{and}\quad\sum_{j=0}^{\infty}\hat{X}_{j}(n)<\infty.$ (32)
The transition probabilities of $\hat{X}(n)$ are given by adapting (4) and (6). Thus, for every $j,n\in\mathbb{N}$ we have that
$\mathbb{P}\left\{\hat{X}_{j}(n+1)\,\big|\,\hat{X}(n)\right\}=\mathbb{P}\left\{\hat{X}_{j}(n+1)\,\big|\,\hat{X}_{j-1}(n),\ \hat{X}_{j}(n)\right\}$ (33)
(with the convention that $\hat{X}_{-1}(n)=0$ for all $n\in\mathbb{N}$) and (compare with (4))
$\left.\begin{array}[]{l}\mathbb{P}\left\{\hat{X}_{j}(n+1)=1\,\big|\,\hat{X}_{j-1}(n)=\hat{X}_{j}(n)=0\right\}=0,\\ \mathbb{P}\left\{\hat{X}_{j}(n+1)=1\,\big|\,\hat{X}_{j-1}(n)=1,\ \hat{X}_{j}(n)=0\right\}=\alpha,\\ \mathbb{P}\left\{\hat{X}_{j}(n+1)=1\,\big|\,\hat{X}_{j-1}(n)=0,\ \hat{X}_{j}(n)=1\right\}=\beta,\\ \mathbb{P}\left\{\hat{X}_{j}(n+1)=1\,\big|\,\hat{X}_{j-1}(n)=\hat{X}_{j}(n)=1\right\}=\gamma.\end{array}\right\}$ (34)
As before, $\alpha$, $\beta$, and $\gamma$ are given constants satisfying (5). To complete the description of the law of $\hat{X}(n)$ we pose that the (Bernoulli) random variables $\hat{X}_{0}(n+1),\hat{X}_{1}(n+1),\hat{X}_{2}(n+1),\ldots$ are conditionally independent given $\hat{X}(n)$, for every $n\in\mathbb{N}$ (this assumption, together with (33), is equivalent to (6)).
The boundary conditions (2) correspond to the initial condition
$\hat{X}_{0}(0)=1\qquad\text{and}\qquad\hat{X}_{j}(0)=0\quad\text{for all }j\geq 1.$ (35)
Finally, let us mention that the valences
$\hat{Y}(n)=\boldsymbol{v}\left(\hat{X}(n)\right)=\sum_{j=0}^{\infty}\hat{X}_{j}(n),\qquad n\in\mathbb{N},$ (36)
correspond to the variables $Y_{n}$, $n\in\mathbb{N}$, introduced in the next section.
3 The random variables $Y_{n}$, $n\geq 0$
It is somehow natural to introduce the random variables
$Y_{n}=\sum_{j=0}^{n}X_{j}^{n},\qquad n\geq 0.$ (37)
Actually, regarding the forest fire, the limiting behavior, as $n\to\infty$, of $Y_{n}$ plays a dominant role in the fate of the forest.
Notice that
$Y_{0}=1\qquad\text{and}\qquad Y_{n}\in\{0,1,\ldots,n+1\}.$ (38)
Also, by the defining properties of the random field $X_{j}^{n}$ (recall (7)) there is a subset $\Omega_{0}$ of $\Omega$ with $\mathbb{P}(\Omega_{0})=1$ such that
$\left\{Y_{n}=0\right\}\cap\Omega_{0}\subset\left\{Y_{n+1}=0\right\}\cap\Omega_{0},\qquad n\geq 0,$ (39)
and, more generally,
$\left\{Y_{n}\leq m\right\}\cap\Omega_{0}\subset\left\{Y_{n+1}\leq 2m\right\}\cap\Omega_{0},\qquad n\geq 0,$ (40)
We have
$\mathbb{E}\left[Y_{n}\,\big|\,\mathcal{F}_{n-1}\right]=\sum_{j=0}^{n}\mathbb{E}\left[X_{j}^{n}\,\big|\,\mathcal{F}_{n-1}\right],\qquad n\geq 1,$ (41)
hence, in view of (11),
$\mathbb{E}\left[Y_{n}\,\big|\,\mathcal{F}_{n-1}\right]=\alpha\sum_{j=0}^{n}X_{j-1}^{n-1}+\beta\sum_{j=0}^{n}X_{j}^{n-1}-(\alpha+\beta-\gamma)\sum_{j=0}^{n}X_{j-1}^{n-1}X_{j}^{n-1},$
from which, and the boundary conditions (2), it follows that
$\mathbb{E}\left[Y_{n}\,\big|\,\mathcal{F}_{n-1}\right]=(\alpha+\beta)Y_{n-1}-(\alpha+\beta-\gamma)\sum_{j=1}^{n-1}X_{j-1}^{n-1}X_{j}^{n-1},\qquad n\geq 1.$ (42)
4 The case $\alpha+\beta=\gamma$
Formula (42) suggests that the case
$\alpha+\beta=\gamma$ (43)
will be easier to analyze. In this section it is always assumed that (43) is satisfied.
Under (43), formula (11) becomes
$\mathbb{P}\left\{X_{j}^{n}=1\,\big|\,\mathcal{F}_{n-1}\right\}=\mathbb{E}\left[X_{j}^{n}\,\big|\,\mathcal{F}_{n-1}\right]=\alpha X_{j-1}^{n-1}+\beta X_{j}^{n-1}.$ (44)
Since $\mathcal{F}_{n-2}\subset\mathcal{F}_{n-1}$, by conditioning on $\mathcal{F}_{n-2}$ formula (44) yields
$\mathbb{E}\left[X_{j}^{n}\,\big|\,\mathcal{F}_{n-2}\right]$ $=\alpha\mathbb{E}\left[X_{j-1}^{n-1}\,\big|\,\mathcal{F}_{n-2}\right]+\beta\mathbb{E}\left[X_{j}^{n-1}\,\big|\,\mathcal{F}_{n-2}\right]$
$=\alpha\left[\alpha X_{j-2}^{n-2}+\beta X_{j-1}^{n-2}\right]+\beta\left[\alpha X_{j-1}^{n-2}+\beta X_{j}^{n-2}\right]$ (45)
$=\alpha^{2}X_{j-2}^{n-2}+2\alpha\beta X_{j-1}^{n-2}+\beta^{2}X_{j}^{n-2}$ (46)
and by straightforward induction
$\mathbb{P}\left\{X_{j}^{n}=1\,\big|\,\mathcal{F}_{n-m}\right\}=\mathbb{E}\left[X_{j}^{n}\,\big|\,\mathcal{F}_{n-m}\right]=\sum_{k=0}^{m}\left({m}\atop{k}\right)\alpha^{k}\beta^{m-k}X_{j-k}^{n-m},\quad 0\leq m\leq n$ (47)
(in the case $m=0$, formula (47) is trivially true since $X_{j}^{n}$ is $\mathcal{F}_{n}$-measurable).
In particular, for $m=n$, in view of (2) and the fact that $\mathcal{F}_{0}=\{\varnothing,\Omega\}$, formula (47) yields
$\mathbb{P}\left\{X_{j}^{n}=1\right\}=\mathbb{E}\left[X_{j}^{n}\right]=\sum_{k=0}^{n}\left({n}\atop{k}\right)\alpha^{k}\beta^{n-k}X_{j-k}^{0}=\left({n}\atop{j}\right)\alpha^{j}\beta^{n-j}.$ (48)
Formula (48) remains valid even in the case where $j$ is not between 0 and $n$, since if $j<0$ or $j>n$, then $X_{j}^{n}=X_{j(n-j)}=0$ and at the same time the binomial coefficient $\left({n}\atop{j}\right)$ vanishes.
Now, regardind $Y_{n}$ we notice that, under (43), formula (42) becomes
$\mathbb{E}\left[Y_{n}\,\big|\,\mathcal{F}_{n-1}\right]=(\alpha+\beta)Y_{n-1}.$ (49)
It follows that if we set
$M_{n}=(\alpha+\beta)^{-n}Y_{n},\qquad n\geq 0,$ (50)
then
$\mathbb{E}\left[M_{n}\,\big|\,\mathcal{F}_{n-1}\right]=(\alpha+\beta)^{-n}\mathbb{E}\left[Y_{n}\,\big|\,\mathcal{F}_{n-1}\right]=(\alpha+\beta)^{-(n-1)}Y_{n-1}=M_{n-1},$ (51)
i.e. $M_{n}$ is an $\mathcal{F}_{n}$-martingale with
$\mathbb{E}\left[M_{n}\right]=\mathbb{E}\left[M_{0}\right]=\mathbb{E}\left[Y_{0}\right]=1\qquad\text{for all }n\geq 0$ (52)
and, consequently, in view of (50),
$\mathbb{E}\left[Y_{n}\right]=\mathbb{E}\left[\left(\alpha+\beta\right)^{n}M_{n}\right]=\left(\alpha+\beta\right)^{n}$ (53)
(which also follows from (48)).
Since $M_{n}\geq 0$ we have that [1]
$M_{n}\to M\ \text{a.s.},$ (54)
where $M$ is a nonnegative random variable satisfying
$0\leq\mathbb{E}[M]\leq\mathbb{E}\left[M_{n}\right]=1.$ (55)
Equation (43) implies that $\alpha+\beta\leq 1$. In the case
$\alpha+\beta<1$ (56)
formulas (50), (54), and (53) imply that
$Y_{n}\to 0\ \text{a.s.}\qquad\text{and}\qquad\mathbb{E}\left[Y_{n}\right]\to 0$ (57)
(in other words $Y_{n}$ converges to 0 almost surely and in the $L_{1}$-sense). Also, in view of (53),
$\mathbb{E}\left[\sum_{n=0}^{\infty}Y_{n}\right]=\sum_{n=0}^{\infty}\mathbb{E}\left[Y_{n}\right]=\sum_{n=0}^{\infty}\left(\alpha+\beta\right)^{n}=\frac{1}{1-\alpha-\beta},$ (58)
where the first equality is justified by Tonelli’s Theorem. Therefore,
$\sum_{n=0}^{\infty}Y_{n}<\infty\ \text{a.s.}$ (59)
which suggests that, if $\alpha+\beta=\gamma<1$, the forest will be, essentially, spared.
Remark 3. The random variable $Y_{n}$ takes values in $\mathbb{N}$. Therefore, the fact that, assuming $\alpha+\beta<1$, we have $Y_{n}\to 0$ a.s. implies that there is a subset $\Omega_{0}$ of $\Omega$ with $\mathbb{P}(\Omega_{0})=1$ with the property that for every $\omega\in\Omega_{0}$ there is an $n_{0}=n_{0}(\omega)$ such that $Y_{n}(\omega)=0$ for every $n\geq n_{0}$. In particular, (59) follows from (57). Also, in view of (50), it follows that $M_{n}(\omega)=0$, too, for every $n\geq n_{0}$. Therefore, (54) becomes $M_{n}\to 0$ a.s., i.e. $M=0$ a.s. However, since $\mathbb{E}\left[M_{n}\right]=1$ for all $n\geq 0$, the sequence $M_{n}$, $n\geq 0$, is not uniformly integrable [1] (of course, in view of (57), $Y_{n}$, $n\geq 0$, is uniformly integrable). Finally, let us notice that $Y_{n}\to 0$ a.s. implies
$\lim_{n}\mathbb{P}\left\{X_{j}^{n}=0,\ 0\leq j\leq n\right\}=\lim_{n}\mathbb{P}\left\{Y_{n}=0\right\}=1.$ (60)
$\diamondsuit$
Remark 4. Suppose $\alpha+\beta<\gamma<1$. Then, we can choose $\tilde{\alpha}\geq\alpha$ and $\tilde{\beta}\geq\beta$ so that $\tilde{\alpha}+\tilde{\beta}=\gamma<1$. Thus, by Remark 1 we get that formula (57) (and, consequently, (59) thanks to the previous remark) continues to hold, while (58) becomes an inequality, namely
$\sum_{n=0}^{\infty}\mathbb{E}\left[Y_{n}\right]\leq\frac{1}{1-\tilde{\alpha}-\tilde{\beta}}$.
The same conclusions hold in the case where $\gamma<\alpha+\beta<1$ (since we can choose $\tilde{\gamma}=\alpha+\beta$). $\diamondsuit$
4.1 The case $\alpha+\beta=\gamma=1$
In the case $\alpha+\beta=\gamma=1$ formula (50) becomes
$M_{n}=Y_{n}$ (61)
and, therefore $Y_{n}$ is an $\mathcal{F}_{n}$-martingale with
$\mathbb{E}\left[Y_{n}\right]=\mathbb{E}\left[Y_{0}\right]=1\qquad\text{for all }n\geq 0.$ (62)
Formula (58) now becomes
$\mathbb{E}\left[\sum_{n=0}^{\infty}Y_{n}\right]=\infty$ (63)
(hence, the fire might cause a serious damage to the forest).
Formula (54) is, of course, still valid, namely
$Y_{n}\to Y\text{a.s.},$ (64)
where here $Y$ is a nonnegative random variable taking values in $\mathbb{N}$ (since $Y_{n}$ takes values in $\mathbb{N}$ for all $n\geq 0$) and satisfying
$0\leq\mathbb{E}[Y]\leq\mathbb{E}\big[Y_{n}\big]=1.$ (65)
Theorem 1. If $\gamma=\alpha+\beta=1$ (with $\alpha\beta>0$, as usual) and $Y_{n}$ is as in (37), then,
$Y_{n}\to 0\ \text{a.s.}$ (66)
Proof. From (64) we know that $Y_{n}\to Y$ a.s., i.e. that the almost sure limit of $Y_{n}$ exists. Hence, we only need to show that $Y=0$ a.s.
Consider the events
$A_{n}=\left\{X_{j}^{n}=1,\ 0\leq j\leq n\right\},\qquad n\geq 0$ (67)
(notice that $A_{0}=\Omega$).
With the help of Example 1 and (48) it is not hard to see that
$\mathbb{P}\left(A_{n}\right)>0\qquad\text{for all }n\geq 0.$ (68)
For notational convenience we denote by $\mathbb{P}_{n}$ the conditional probability given $A_{n}$, that is for any event $B$ we have
$\mathbb{P}_{n}(B)=\mathbb{P}(B|A_{n}),\qquad n\geq 0$ (69)
(of course, $\mathbb{P}_{0}=\mathbb{P}$).
Now, let
$x_{n}=\mathbb{P}_{n}\left\{Y=0\right\}=\mathbb{P}\left\{Y=0\,\big|\,A_{n}\right\}=\mathbb{P}\left\{Y=0\,\big|\,X_{j}^{n}=1,\ 0\leq j\leq n\right\},\quad n\geq 0.$ (70)
Then, we have
$x_{0}$ $=\mathbb{P}\left\{Y=0\right\}$
$=\sum_{\epsilon_{0},\epsilon_{1}=0\text{ or }1}\mathbb{P}\left\{Y=0\,\big|\,X_{0}^{1}=\epsilon_{0},\ X_{1}^{1}=\epsilon_{1}\right\}\mathbb{P}\left\{X_{0}^{1}=\epsilon_{0},\ X_{1}^{1}=\epsilon_{1}\right\}$
$=\sum_{\epsilon_{0},\epsilon_{1}=0\text{ or }1}\mathbb{P}\left\{Y=0\,\big|\,X_{0}^{1}=\epsilon_{0},\ X_{1}^{1}=\epsilon_{1}\right\}\mathbb{P}\left\{X_{0}^{1}=\epsilon_{0}\right\}\mathbb{P}\left\{X_{1}^{1}=\epsilon_{1}\right\}$ (71)
(the second equality follows from (6) and the fact that $\mathcal{G}_{0}$ is the trivial $\sigma$-algebra). Notice that the sum in (71) consists of four terms.
The next three formulas follow from (4) and (70).
$\mathbb{P}\left\{Y=0\,\big|\,X_{0}^{1}=0,\ X_{1}^{1}=0\right\}=1,$ (72)
$\mathbb{P}\left\{Y=0\,\big|\,X_{0}^{1}=0,\ X_{1}^{1}=1\right\}=\mathbb{P}\left\{Y=0\,\big|\,X_{0}^{1}=1,\ X_{1}^{1}=0\right\}=x_{0}$ (73)
(both probabilities in (73) are equal to the conditional probability of the event $\{Y=0\}$ given that the fire started at one point, which is exactly the probability denoted by $x_{0}$),
$\mathbb{P}\left\{Y=0\,\big|\,X_{0}^{1}=1,\ X_{1}^{1}=1\right\}=x_{1}.$ (74)
Substituting (72), (73), and (74) in (71) yields
$x_{0}$ $=\mathbb{P}\left\{X_{0}^{1}=0\right\}\mathbb{P}\left\{X_{1}^{1}=0\right\}+x_{0}\mathbb{P}\left\{X_{0}^{1}=0\right\}\mathbb{P}\left\{X_{1}^{1}=1\right\}$
$+x_{0}\mathbb{P}\left\{X_{0}^{1}=1\right\}\mathbb{P}\left\{X_{1}^{1}=0\right\}+x_{1}\mathbb{P}\left\{X_{0}^{1}=1\right\}\mathbb{P}\left\{X_{1}^{1}=1\right\}.$ (75)
The above probabilities can be computed by (48) (and the assumption that $\alpha+\beta=1$). We have
$\mathbb{P}\left\{X_{0}^{1}=0\right\}=1-\beta=\alpha,\qquad\mathbb{P}\left\{X_{0}^{1}=1\right\}=\beta,$ (76)
and
$\mathbb{P}\left\{X_{1}^{1}=0\right\}=1-\alpha=\beta,\qquad\mathbb{P}\left\{X_{0}^{1}=1\right\}=\alpha.$ (77)
Hence, by using (76) and (77) in (75) we get
$x_{0}=\alpha\beta+\alpha^{2}x_{0}+\beta^{2}x_{0}+\alpha\beta x_{1}.$ (78)
Using the fact that $\alpha^{2}+\beta^{2}=1-2\alpha\beta$, formula (78) yields
$x_{0}=\frac{1+x_{1}}{2}>\frac{1}{2}.$ (79)
Next, we will look for a recursive formula for $x_{n}$.
$x_{n}$ $=\mathbb{P}_{n}\left\{Y=0\right\}$
$=\sum_{\epsilon_{j}=0\text{ or }1\atop 0\leq j\leq n+1}\left[\,\mathbb{P}_{n}\left\{Y=0\,\big|\,X_{j}^{n+1}=\epsilon_{j},\ 0\leq j\leq n+1\right\}\right$.
$\qquad\qquad\ \ \left.\times\,\mathbb{P}_{n}\left\{X_{j}^{n+1}=\epsilon_{j},\ 0\leq j\leq n+1\right\}\right]$ (80)
From (67) and (3) we have that $A_{n}\in\mathcal{G}_{n}$. Hence, in view of (6) and (69), the events $\{X_{j}^{n+1}=\epsilon_{j}\}$, $0\leq j\leq n+1$, are independent with respect to the probability $\mathbb{P}_{n}$ and, consequently, formula (80) becomes
$x_{n}$ $=\mathbb{P}_{n}\left\{Y=0\right\}$
$=\sum_{\epsilon_{j}=0\text{ or }1\atop 0\leq j\leq n+1}\left[\,\mathbb{P}_{n}\left\{Y=0\,\big|\,X_{j}^{n+1}=\epsilon_{j},\ 0\leq j\leq n+1\right\}\right$.
$\qquad\qquad\ \ \times\,\prod_{j=0}^{n+1}\mathbb{P}_{n}\left\{X_{j}^{n+1}=\epsilon_{j}\right\}\big].$ (81)
Now, in view of (4), the assumption that $\gamma=1$ implies (recall (67) and (69))
$\mathbb{P}_{n}\left\{X_{j}^{n+1}=0\right\}=0\,\,\Longleftrightarrow\,\,\mathbb{P}_{n}\left\{X_{j}^{n+1}=1\right\}=1\qquad\text{for every }j\in\{1,2,\ldots,n\},$ (82)
thus, by using (82) and (81) we obtain
$x_{n}$ $=\mathbb{P}_{n}\left\{Y=0\right\}$
$=\sum_{\epsilon_{0},\epsilon_{1}=0\text{ or }1}\left[\,\mathbb{P}_{n}\left\{Y=0\,\big|\,X_{0}^{n+1}=\epsilon_{0},\ X_{n+1}^{n+1}=\epsilon_{1},\ X_{j}^{n+1}=1,\ 1\leq j\leq n\right\}\right$.
$\qquad\qquad\qquad\ \ \left.\times\,\mathbb{P}_{n}\left\{X_{0}^{n+1}=\epsilon_{0}\right\}\mathbb{P}_{n}\left\{X_{n+1}^{n+1}=\epsilon_{1}\right\}\right]$
$=\sum_{\epsilon_{0},\epsilon_{1}=0\text{ or }1}\left[\,\mathbb{P}\left\{Y=0\,\big|\,X_{0}^{n+1}=\epsilon_{0},\ X_{n+1}^{n+1}=\epsilon_{1},\ X_{j}^{n+1}=1,\ 1\leq j\leq n\right\}\right$.
$\qquad\qquad\qquad\ \ \left.\times\,\mathbb{P}_{n}\left\{X_{0}^{n+1}=\epsilon_{0}\right\}\mathbb{P}_{n}\left\{X_{n+1}^{n+1}=\epsilon_{1}\right\}\right],$ (83)
where the last equality (i.e, to use $\mathbb{P}$ instead of $\mathbb{P}_{n}$ in the conditional probability) follows from the first equality of (6) and the fact that (in view of (3))
$\left\{X_{0}^{n+1}=\epsilon_{0},\ X_{n+1}^{n+1}=\epsilon_{1},\ X_{j}^{n+1}=1,\ 1\leq j\leq n\right\}\in\mathcal{G}_{n+1}.$
In view of (70), we have for $n\geq 1$
$\mathbb{P}\left\{Y=0\,\big|\,X_{0}^{n+1}=0,\ X_{n+1}^{n+1}=0,\ X_{j}^{n+1}=1,\ 1\leq j\leq n\right\}=x_{n-1},$ (84)
$\mathbb{P}\left\{Y=0\,\big|\,X_{0}^{n+1}=0,\ X_{n+1}^{n+1}=1,\ X_{j}^{n+1}=1,\ 1\leq j\leq n\right\}=x_{n},$ (85)
$\mathbb{P}\left\{Y=0\,\big|\,X_{0}^{n+1}=1,\ X_{n+1}^{n+1}=0,\ X_{j}^{n+1}=1,\ 1\leq j\leq n\right\}=x_{n},$ (86)
and
$\mathbb{P}\left\{Y=0\,\big|\,X_{0}^{n+1}=1,\ X_{n+1}^{n+1}=1,\ X_{j}^{n+1}=1,\ 1\leq j\leq n\right\}=x_{n+1}.$ (87)
Substituting (84), (85), (86), and (87) in (83) yields
$x_{n}$ $=x_{n-1}\mathbb{P}_{n}\left\{X_{0}^{n+1}=0\right\}\mathbb{P}_{n}\left\{X_{n+1}^{n+1}=0\right\}+x_{n}\mathbb{P}_{n}\left\{X_{0}^{n+1}=0\right\}\mathbb{P}_{n}\left\{X_{n+1}^{n+1}=1\right\}$
$+x_{n}\mathbb{P}_{n}\left\{X_{0}^{n+1}=1\right\}\mathbb{P}_{n}\left\{X_{n+1}^{n+1}=0\right\}+x_{n+1}\mathbb{P}_{n}\left\{X_{0}^{n+1}=1\right\}\mathbb{P}_{n}\left\{X_{n+1}^{n+1}=1\right\}.$ (88)
Finally, we calculate the probabilities that appear in (88). In view of (67), (69),(6), (4), and (2), we get
$\mathbb{P}_{n}\left\{X_{0}^{n+1}=0\right\}$ $=\mathbb{P}\left\{X_{0}^{n+1}=0\,\big|\,X_{j}^{n}=1,\ 0\leq j\leq n\right\}$
$=\mathbb{P}\left\{X_{0}^{n+1}=0\,\big|\,X_{0}^{n}=1\right\}=1-\beta=\alpha$ (89)
and
$\mathbb{P}_{n}\left\{X_{n+1}^{n+1}=0\right\}$ $=\mathbb{P}\left\{X_{n+1}^{n+1}=0\,\big|\,X_{j}^{n}=1,\ 0\leq j\leq n\right\}$
$=\mathbb{P}\left\{X_{n+1}^{n+1}=0\,\big|\,X_{n}^{n}=1\right\}=1-\alpha=\beta,$ (90)
while from (89) and (90) we get immediately that
$\mathbb{P}_{n}\left\{X_{0}^{n+1}=1\right\}=\beta\qquad\text{and}\qquad\mathbb{P}_{n}\left\{X_{n+1}^{n+1}=1\right\}=\alpha.$ (91)
Therefore, by substituting (89), (90), and (91), in (88) we obtain
$x_{n}=\alpha\beta x_{n-1}+\alpha^{2}x_{n}+\beta^{2}x_{n}+\alpha\beta x_{n+1},$ (92)
which implies (since $\alpha^{2}+\beta^{2}=1-2\alpha\beta$)
$x_{n}=\frac{x_{n-1}+x_{n+1}}{2}\qquad n\geq 1.$ (93)
Equation (93) is a very simple second-order linear difference equation with constant coefficients. Its general solution is
$x_{n}=c_{1}+c_{2}n,\qquad n\geq 0,$
where $c_{1},c_{2}$ are constants. However, in our case $x_{n}$ is a probability. Hence, we must have $c_{2}=0$ and, consequently, $x_{n}$ is constant, i.e.
$x_{n}=x_{0}\qquad n\geq 1.$ (94)
In particular, $x_{1}=x_{0}$, hence (79) becomes
$x_{0}=\frac{1+x_{0}}{2}$
which implies that $x_{0}=1$, i.e. (in view of (70)) $\mathbb{P}\{Y=0\}=1$. $\blacksquare$
The symbol $\blacksquare$ indicates the end of a proof.
Actually, in the proof of Theorem 1 we have shown that
$x_{n}=\mathbb{P}\left\{Y=0\,\big|\,X_{j}^{n}=1,\ 0\leq j\leq n\right\}=1\qquad\text{for all }n\geq 0,$ (95)
which is stronger than $\mathbb{P}\{Y=0\}=1$.
Remark 5. From Theorem 1 we see that $\mathbb{E}[Y]=0\neq 1=\mathbb{E}[Y_{n}]$. Hence, in the case $\gamma=\alpha+\beta=1$ the sequence $Y_{n}$, $n\geq 0$, is not uniformly integrable [1]. $\diamondsuit$
Remark 6. Suppose $\alpha+\beta=1$ and $\gamma<1$. Then, from formula (42) we get that
$\mathbb{E}\left[Y_{n}\,\big|\,\mathcal{F}_{n-1}\right]\leq Y_{n-1},$
i.e. $Y_{n}$ is a (nonnegative) $\mathcal{F}_{n}$-supermartingale. Therefore, (64) continues to hold [1]. Consequently, in view of Remark 1, we get that Theorem 1 and, furthermore, (95) are still valid. $\diamondsuit$
Regarding the fate of the forest, the results of this subsection lead us to the conclusion that $\alpha+\beta=\gamma=1$ is a borderline case. On one hand formula (63) suggests that many trees will be burnt, while on the other hand formula (95) tells us that eventually the fire will (eventually) die out.
4.2 The growth of $\mathbb{E}\left[Y_{n}^{2}\right]$
In this subsection we focus on $Y_{n}^{2}$. Actually, our main interest is the behavior of $\mathbb{E}\left[Y_{n}^{2}\right]$ as $n\to\infty$.
From the defining formula (37) of $Y_{n}$ we get that
$Y_{n}^{2}=\left(\sum_{j=0}^{n}X_{j}^{n}\right)^{2}$ $=\sum_{j=0}^{n}\left(X_{j}^{n}\right)^{2}+2\sum_{j=0}^{n}\sum_{k=j+1}^{n}X_{j}^{n}X_{k}^{n}$
$=\sum_{j=0}^{n}X_{j}^{n}+2\sum_{j=0}^{n}\sum_{k=j+1}^{n}X_{j}^{n}X_{k}^{n}$
$=Y_{n}+2\sum_{j=0}^{n}\sum_{k=j+1}^{n}X_{j}^{n}X_{k}^{n},\qquad\qquad n\geq 0.$ (96)
Hence,
$\mathbb{E}\left[Y_{n}^{2}\right]=\mathbb{E}\left[Y_{n}\right]+2\sum_{j=0}^{n}\sum_{k=j+1}^{n}\mathbb{E}\left[X_{j}^{n}X_{k}^{n}\right],\qquad n\geq 0.$ (97)
Throughout this section we assume (43), i.e. $\gamma=\alpha+\beta$. Therefore, (53) holds and, consequently, (97) becomes
$\mathbb{E}\left[Y_{n}^{2}\right]=\gamma^{n}+2\sum_{j=0}^{n}\sum_{k=j+1}^{n}\mathbb{E}\left[X_{j}^{n}X_{k}^{n}\right],\qquad n\geq 0.$ (98)
Let us set
$R_{0}=0,\qquad R_{n}=\sum_{j=0}^{n}\sum_{k=j+1}^{n}\mathbb{E}\left[X_{j}^{n}X_{k}^{n}\right],\qquad n\geq 1,$ (99)
so that (98) can be written as
$\mathbb{E}\left[Y_{n}^{2}\right]=\gamma^{n}+2R_{n},\qquad n\geq 0.$ (100)
Now, under (43) formula (13) becomes
$\mathbb{E}\left[X_{j}^{n}X_{k}^{n}\,\big|\,\mathcal{F}_{n-1}\right]=\left(\alpha X_{j-1}^{n-1}+\beta X_{j}^{n-1}\right)\left(\alpha X_{k-1}^{n-1}+\beta X_{k}^{n-1}\right)$
$=\alpha^{2}X_{j-1}^{n-1}X_{k-1}^{n-1}+\alpha\beta X_{j-1}^{n-1}X_{k}^{n-1}+\alpha\beta X_{j}^{n-1}X_{k-1}^{n-1}+\beta^{2}X_{j}^{n-1}X_{k}^{n-1},$ (101)
hence, by taking expectations we obtain
$\mathbb{E}\left[X_{j}^{n}X_{k}^{n}\right]$ $=\alpha^{2}\mathbb{E}\left[X_{j-1}^{n-1}X_{k-1}^{n-1}\right]+\alpha\beta\mathbb{E}\left[X_{j-1}^{n-1}X_{k}^{n-1}\right]$
$+\alpha\beta\mathbb{E}\left[X_{j}^{n-1}X_{k-1}^{n-1}\right]+\beta^{2}\mathbb{E}\left[X_{j}^{n-1}X_{k}^{n-1}\right].$ (102)
Using (102) in (99) yields
$R_{n}$ $=\alpha^{2}\sum_{j=0}^{n}\sum_{k=j+1}^{n}\mathbb{E}\left[X_{j-1}^{n-1}X_{k-1}^{n-1}\right]+\alpha\beta\sum_{j=0}^{n}\sum_{k=j+1}^{n}\mathbb{E}\left[X_{j-1}^{n-1}X_{k}^{n-1}\right]$
$+\alpha\beta\sum_{j=0}^{n}\sum_{k=j+1}^{n}\mathbb{E}\left[X_{j}^{n-1}X_{k-1}^{n-1}\right]+\beta^{2}\sum_{j=0}^{n}\sum_{k=j+1}^{n}\mathbb{E}\left[X_{j}^{n-1}X_{k}^{n-1}\right].$ (103)
In view of the boundary conditions (2), formula (103) becomes
$R_{n}$ $=\alpha^{2}\sum_{j=0}^{n-1}\sum_{k=j+1}^{n-1}\mathbb{E}\left[X_{j}^{n-1}X_{k}^{n-1}\right]+\alpha\beta\sum_{j=0}^{n-1}\sum_{k=j+2}^{n-1}\mathbb{E}\left[X_{j}^{n-1}X_{k}^{n-1}\right]$
$+\alpha\beta\sum_{j=0}^{n-1}\sum_{k=j}^{n-1}\mathbb{E}\left[X_{j}^{n-1}X_{k}^{n-1}\right]+\beta^{2}\sum_{j=0}^{n-1}\sum_{k=j+1}^{n-1}\mathbb{E}\left[X_{j}^{n-1}X_{k}^{n-1}\right]$
$=(\alpha^{2}+\beta^{2})R_{n-1}$
$+\alpha\beta\sum_{j=0}^{n-1}\sum_{k=j+2}^{n-1}\mathbb{E}\left[X_{j}^{n-1}X_{k}^{n-1}\right]+\alpha\beta\sum_{j=0}^{n-1}\sum_{k=j}^{n-1}\mathbb{E}\left[X_{j}^{n-1}X_{k}^{n-1}\right]$
$=(\alpha^{2}+\beta^{2})R_{n-1}$
$+\alpha\beta\left(R_{n-1}-\sum_{j=0}^{n-1}\mathbb{E}\left[X_{j}^{n-1}X_{j+1}^{n-1}\right]\right)+\alpha\beta\left(R_{n-1}+\sum_{j=0}^{n-1}\mathbb{E}\left[X_{j}^{n-1}X_{j}^{n-1}\right]\right)$
$=(\alpha+\beta)^{2}R_{n-1}-\alpha\beta\sum_{j=0}^{n-2}\mathbb{E}\left[X_{j}^{n-1}X_{j+1}^{n-1}\right]+\alpha\beta\sum_{j=0}^{n-1}\mathbb{E}\left[X_{j}^{n-1}\right]$
$=(\alpha+\beta)^{2}R_{n-1}-\alpha\beta\sum_{j=0}^{n-2}\mathbb{E}\left[X_{j}^{n-1}X_{j+1}^{n-1}\right]+\alpha\beta\mathbb{E}\left[Y_{n-1}\right],$ (104)
hence, in view of (43) and (53),
$R_{n}=\gamma^{2}R_{n-1}+\alpha\beta\gamma^{n-1}-\alpha\beta\sum_{j=0}^{n-2}\mathbb{E}\left[X_{j}^{n-1}X_{j+1}^{n-1}\right],\qquad n\geq 1.$ (105)
One very simple implication of formula (105) is the inequality
$R_{n}\leq\gamma^{2}R_{n-1}+\alpha\beta\gamma^{n-1},\qquad n\geq 1.$ (106)
Case $\gamma=\alpha+\beta<1$. In this case (106) implies
$\gamma^{-2n}R_{n}-\gamma^{-2(n-1)}R_{n-1}\leq\alpha\beta\gamma^{-(n+1)}\qquad n\geq 1$
hence,
$\sum_{n=1}^{N}\left(\gamma^{-2n}R_{n}-\gamma^{-2(n-1)}R_{n-1}\right)\leq\alpha\beta\sum_{n=1}^{N}\gamma^{-(n+1)},$
i.e.
$\gamma^{-2N}R_{N}\leq\alpha\beta\sum_{n=1}^{N}\gamma^{-(n+1)}=\frac{\alpha\beta}{\gamma}\frac{\gamma^{-N}-1}{1-\gamma},\qquad N\geq 1,$
from which it follows that
$R_{n}\leq\alpha\beta\frac{1-\gamma^{n}}{1-\gamma}\gamma^{n-1},\qquad n\geq 1.$ (107)
Finally, using (107) in (100) yields
$\mathbb{E}\left[Y_{n}^{2}\right]\leq\left(\gamma+2\alpha\beta\frac{1-\gamma^{n}}{1-\gamma}\right)\gamma^{n-1},\qquad n\geq 1,$ (108)
hence if $\gamma=\alpha+\beta<1$ (more generally, if $\gamma\leq\alpha+\beta<1$), then, as $n\to\infty$, $\mathbb{E}\left[Y_{n}^{2}\right]$ aproaches 0 exponentially fast. In particular, $Y_{n}\to 0$ in the $L_{2}$-sense.
Case $\gamma=\alpha+\beta=1$. Here the fact that (recall (51) and (61)) $Y_{n}$ is an $\mathcal{F}_{n}$-martingale implies [1] that $Y_{n}^{2}$ is an $\mathcal{F}_{n}$-submartingale. In particular,
$\mathbb{E}\left[Y_{n}^{2}\right]\geq\mathbb{E}\left[Y_{n-1}^{2}\right],\qquad n\geq 1.$ (109)
Then, the fact that the sequence $Y_{n}$, $n\geq 0$, is not uniformly integrable (Remark 5) implies that
$\lim_{n}\mathbb{E}\left[Y_{n}^{2}\right]=\infty,$ (110)
since, if $\mathbb{E}\left[Y_{n}^{2}\right]$ were bounded, then $Y_{n}$, $n\geq 0$, would have been uniformly integrable [1].
We can easily get an upper estimate of the growth of $\mathbb{E}\left[Y_{n}^{2}\right]$. From (100) (since here $\gamma=1$) we get that
$\mathbb{E}\left[Y_{n}^{2}\right]=1+2R_{n},\qquad n\geq 0.$ (111)
Now, formula (105) becomes
$R_{n}=R_{n-1}+\alpha\beta-\alpha\beta\sum_{j=0}^{n-2}\mathbb{E}\left[X_{j}^{n-1}X_{j+1}^{n-1}\right],\qquad n\geq 1,$ (112)
while inequality (106) becomes
$R_{n}\leq R_{n-1}+\alpha\beta,\qquad n\geq 1.$ (113)
Thus, in view of (111), the inequality (113) implies immediately the upper estimate
$\mathbb{E}\left[Y_{n}^{2}\right]\leq 1+2\alpha\beta n,\qquad n\geq 0,$ (114)
hence $\mathbb{E}\left[Y_{n}^{2}\right]$ grows at most linearly.
5 The case $\alpha+\beta>1$
In the case where $\alpha+\beta>1$ we have not established the existence of the almost sure limit of $Y_{n}$. However, if we set
$W_{n}={\bf 1}_{\{Y_{n}=0\}},\qquad n\geq 0,$ (115)
then formula (39) implies
$W_{n}\leq W_{n+1}\text{ a.s.},\qquad n\geq 0,$ (116)
hence,
$W_{n}\to W\text{ a.s.},$ (117)
where $W$ is a Bernoulli random variable. Thus, by the Monotone Convergence Theorem we obtain that (since $\mathbb{P}\left\{Y_{n}=0\right\}=\mathbb{E}\left[W_{n}\right]$)
$\lim_{n}\mathbb{P}\left\{Y_{n}=0\right\}=\lim_{n}\mathbb{E}\left[W_{n}\right]=\mathbb{E}[W].$ (118)
If $\alpha=1$, then the definition of the random field $X_{j}^{n}$ implies that $X_{n}^{n}=1$ a.s. and, consequently, $\mathbb{P}\left\{Y_{n}=0\right\}=0$ for all $n\geq 0$, hence $W=0$ a.s. Likewise, if $\beta=1$, then $X_{0}^{n}=1$ a.s. hence, again $\mathbb{P}\left\{Y_{n}=0\right\}=0$ for all $n\geq 0$ and $W=0$ a.s. For this reason, from now on we will assume that
$\alpha,\beta\neq 1.$ (119)
Then, from the definition of the random field $X_{j}^{n}$ it is not hard to see that
$\mathbb{P}\left\{Y_{n}=0\right\}<\mathbb{P}\left\{Y_{n+1}=0\right\},\qquad n\geq 0.$ (120)
5.1 The case $\alpha+\beta>1$ and $\gamma=1$
Before giving the first result of this subsection let us notice that the assumption that $\alpha+\beta>1$ implies that $(1-\alpha)(1-\beta)<\alpha\beta$ and, hence,
$\frac{(1-\alpha)(1-\beta)}{\alpha\beta}<1.$ (121)
The first result of the subsection is a lower bound of the limit $\lim_{n}\mathbb{P}\left\{Y_{n}=0\right\}$.
Theorem 2. If $\alpha+\beta>1$ and $\gamma=1$ (with $\alpha,\beta\neq 1$), while $Y_{n}$ is as in (37), then
$\lim_{n}\mathbb{P}\left\{Y_{n}=0\right\}\geq\frac{(1-\alpha)(1-\beta)}{\alpha\beta}.$ (122)
Proof. Imitating the proof of Theorem 1 we set (recall (67), (69), and (70))
$x_{n}=\mathbb{E}_{n}[W]=\mathbb{E}\left[W\,\big|\,A_{n}\right]=\mathbb{E}\left[W\,\big|\,X_{j}^{n}=1,\ 0\leq j\leq n\right],\qquad n\geq 0$ (123)
($\mathbb{E}_{n}$ is the expectation corresponding to the probability $\mathbb{P}_{n}$).
It is clear from (123) that
$x_{n}\geq x_{n+1},\qquad n\geq 0.$ (124)
In view of (118) (and the fact that $\mathbb{E}_{0}=\mathbb{E}$),
$x_{0}=\mathbb{E}[W]=\lim_{n}\mathbb{P}\left\{Y_{n}=0\right\}.$ (125)
Also, in view of (120) and (123) we have that
$x_{0}>\mathbb{P}\left\{Y_{1}=0\right\}$ $=\mathbb{P}\left\{X_{0}^{1}=X_{1}^{1}=0\right\}$
$=\mathbb{P}\left\{X_{0}^{1}=0\right\}\mathbb{P}\left\{X_{1}^{1}=0\right\}=(1-\beta)(1-\alpha)$ (126)
(the second equality follows from (6) and the fact that $\mathcal{G}_{0}$ is the trivial $\sigma$-algebra; the third equality follows from (48)).
At this point we notice that (with $\gamma=1$) all formulas in the proof of Theorem 1 remain valid as long as they do not depend on the relation $\alpha+\beta=1$. In particular, formula (75) yields
$x_{0}=(1-\beta)(1-\alpha)+[(1-\beta)\alpha+\beta(1-\alpha)]x_{0}+\alpha\beta x_{1},$
or, equivalently,
$(1-\alpha)(1-\beta)+\alpha\beta]x_{0}=(1-\alpha)(1-\beta)+\alpha\beta x_{1},$ (127)
while formula (88) yields
$x_{n}=(1-\beta)(1-\alpha)x_{n-1}+[(1-\beta)\alpha+\beta(1-\alpha)]x_{n}+\alpha\beta x_{n+1},\qquad n\geq 1,$
or, equivalently,
$(1-\alpha)(1-\beta)+\alpha\beta]x_{n}=(1-\alpha)(1-\beta)x_{n-1}+\alpha\beta x_{n+1},\qquad n\geq 1.$ (128)
Equation (128) is a very simple second-order linear difference equation with constant coefficients. Its general solution is
$x_{n}=c_{1}+c_{2}\left[\frac{(1-\alpha)(1-\beta)}{\alpha\beta}\right]^{n},\qquad n\geq 0,$ (129)
where $c_{1},c_{2}$ are constants.
Using (129) in (127) yields
$(1-\alpha)(1-\beta)+\alpha\beta](c_{1}+c_{2})=(1-\alpha)(1-\beta)+\alpha\beta\left[c_{1}+c_{2}\frac{(1-\alpha)(1-\beta)}{\alpha\beta}\right],$
which implies that
$c_{2}=\frac{(1-\alpha)(1-\beta)}{\alpha\beta}(1-c_{1}).$ (130)
Thus, (129) becomes
$x_{n}=c_{1}+(1-c_{1})\left[\frac{(1-\alpha)(1-\beta)}{\alpha\beta}\right]^{n+1},\qquad n\geq 0,$ (131)
and the fact that $x_{n}$ is a decreasing sequence of probabilities implies (in view of (121)) that
$0\leq c_{1}\leq 1.$ (132)
Finally, for $n=0$ formula (131) yields
$\lim_{n}\mathbb{P}\left\{Y_{n}=0\right\}=x_{0}=c_{1}+(1-c_{1})\frac{(1-\alpha)(1-\beta)}{\alpha\beta},$ (133)
from which (122) follows. $\blacksquare$
Remark 7. The inequality (122) gives a lower bound for $\lim_{n}\mathbb{P}\left\{Y_{n}=0\right\}$. In the extreme case where $\alpha=1$ or $\beta=1$ we have seen that $\mathbb{P}\left\{Y_{n}=0\right\}=0$ for all $n\geq 0$, hence (122) becomes equality. Also, in the case $\alpha+\beta=1$ the right-hand side of (122) is equal to $1$, hence, again (122) becomes equality. It is tempting to conjecture that (122) becomes equality whenever $\alpha+\beta>1$ and $\gamma=1$. $\diamondsuit$
The next theorem gives upper and lower bounds for the expectation $\mathbb{E}\left[Y_{n}\right]$.
Theorem 3. If $\alpha+\beta>1$ and $\gamma=1$ (with $\alpha,\beta\neq 1$), while $Y_{n}$ is as in (37), then:
(i) We have the upper bound
$\mathbb{E}\left[Y_{n}\right]\geq 1+(\alpha+\beta-1)n,\qquad n\geq 0.$ (134)
(ii) For any $\varepsilon>0$ there is a $n_{0}=n_{0}(\varepsilon)$ such that
$\mathbb{E}\left[Y_{n}\right]\leq\left(\frac{\alpha+\beta-1}{\alpha\beta}+\varepsilon\right)n,\qquad n\geq n_{0}$ (135)
(notice that $(\alpha+\beta-1)/(\alpha\beta)<1$), hence
$\limsup_{n}\frac{\mathbb{E}\left[Y_{n}\right]}{n}\leq\frac{\alpha+\beta-1}{\alpha\beta}.$ (136)
Proof. (i) For $\gamma=1$ formula (42) becomes
$\mathbb{E}\left[Y_{n}\,\big|\,\mathcal{F}_{n-1}\right]=Y_{n-1}+(\alpha+\beta-1)\left(Y_{n-1}-\sum_{j=1}^{n-1}X_{j-1}^{n-1}X_{j}^{n-1}\right),\quad n\geq 1.$ (137)
Now, the fact that
$Y_{n-1}=\sum_{j=0}^{n-1}X_{j}^{n-1}=\sum_{j=1}^{n}X_{j-1}^{n-1}$
implies (since the $X_{j}^{n-1}$’s are Bernoulli random variables)
$Y_{n-1}\geq 1+\sum_{j=1}^{n-1}X_{j-1}^{n-1}X_{j}^{n-1}.$ (138)
Thus, (137) implies
$\mathbb{E}\left[Y_{n}\,\big|\,\mathcal{F}_{n-1}\right]\geq Y_{n-1}+\alpha+\beta-1,\qquad n\geq 1,$ (139)
and, therefore, by taking expectations we get
$\mathbb{E}\left[Y_{n}\right]\geq\mathbb{E}\left[Y_{n-1}\right]+\alpha+\beta-1,\qquad n\geq 1$ (140)
(also, since $\alpha+\beta>1$, $Y_{n}$ is an $\mathcal{F}_{n}$-submartingale), and (134) follows immediately from (140) and the fact that $Y_{0}=1$.
(ii) From (122) we get that, given $\varepsilon_{1}>0$ there is a $n_{1}=n_{1}(\varepsilon_{1})$ such that
$\mathbb{P}\left\{Y_{n}=0\right\}\geq\frac{(1-\alpha)(1-\beta)}{\alpha\beta}-\varepsilon_{1},\qquad n\geq n_{1}.$ (141)
Now, since $Y_{n}\in\{0,1,\ldots,n+1\}$,
$\mathbb{E}\left[Y_{n}\right]\leq(n+1)\mathbb{P}\left\{Y_{n}\geq 1\right\}=(n+1)\left(1-\mathbb{P}\left\{Y_{n}=0\right\}\right),$ (142)
hence, in view of (141),
$\mathbb{E}\left[Y_{n}\right]$ $\leq(n+1)\left(1-\frac{(1-\alpha)(1-\beta)}{\alpha\beta}+\varepsilon_{1}\right)$
$=\left(\frac{\alpha+\beta-1}{\alpha\beta}+\varepsilon_{1}\right)(n+1),\qquad n\geq n_{1},$
from which we easily obtain (135). $\blacksquare$
From the inequality (134) we can easily get an upper bound for the probability $\mathbb{P}\left\{Y_{n}=0\right\}$ and its limit $\lim_{n}\mathbb{P}\left\{Y_{n}=0\right\}$.
Corollary 1. If $\alpha+\beta>1$ and $\gamma=1$ (with $\alpha,\beta\neq 1$), while $Y_{n}$ is as in (37), then
$\mathbb{P}\left\{Y_{n}=0\right\}\leq\frac{n}{n+1}(2-\alpha-\beta),\qquad n\geq 0,$ (143)
hence
$\lim_{n}\mathbb{P}\left\{Y_{n}=0\right\}\leq 2-\alpha-\beta.$ (144)
Proof. Using (134) in (144) yields
$1+(\alpha+\beta-1)n\leq(n+1)\left(1-\mathbb{P}\left\{Y_{n}=0\right\}\right),$ (145)
from which (143) follows. $\blacksquare$
Formulas (134) and (144) suggest that, in the case where $\alpha+\beta>1$ and $\gamma=1$ the forest will be considerably damaged by the fire.
Remark 8. By Remark 1 formulas (122) and (135) remain valid in the case $\gamma<1$. $\diamondsuit$
5.2 The random variables $T_{n}$, $n\geq 1$
We set
$T_{n}=\frac{Y_{n}}{n},\qquad n\geq 1.$ (146)
Recall that $Y_{n}\in\{0,1,\ldots,n+1\}$. Hence,
$0\leq T_{n}\leq\frac{n+1}{n}\leq 2,\qquad n\geq 1$ (147)
(thus, for example, $\mathbb{E}\left[T_{n}^{2}\right]\leq 4$ for all $n\geq 1$).
Formula (147) implies [1] that the sequence $\{T_{n}\}_{n\geq 1}$ is uniformly integrable and, also, that the sequence of the distribution functions of $T_{n}$, $n\geq 1$, is tight.
Let us observe that from formulas (134) and (136) it follows immediately that
$\alpha+\beta-1\leq\liminf_{n}\mathbb{E}\left[T_{n}\right]\leq\limsup_{n}\mathbb{E}\left[T_{n}\right]\leq\frac{\alpha+\beta-1}{\alpha\beta}.$ (148)
Open Question. Does $T_{n}$ converge, at least in distribution?
Notice that if $T_{n}$ converges in distribution to a random variable $T$, then we must also have that $\lim_{n}\mathbb{E}\left[T_{n}\right]=\mathbb{E}[T]$. This follows from the fact that the distributional convergence of $T_{n}$ implies [1] that there are random variables $\tilde{T}_{n}$, $n\geq 1$, such that $\tilde{T}_{n}$ and $T_{n}$ have the same distribution (for every $n\geq 1$) and $\tilde{T}_{n}\to\tilde{T}$ a.s., where $\tilde{T}$ and $T$ have the same distribution. Furthermore, (147) implies that $\{\tilde{T}_{n}\}_{n\geq 1}$, too, is uniformly integrable, hence $\lim_{n}\mathbb{E}\left[T_{n}\right]=\lim_{n}\mathbb{E}[\tilde{T}_{n}]=\mathbb{E}[\tilde{T}]=\mathbb{E}[T]$.
Finally, let us introduce some random variables which may be helpful in the study of the random field $X^{n}_{j}$. We set
$L_{0}=\inf\left\{n\geq 0:X^{n}_{0}=0\right\}$ (149)
and
$L_{m}=\inf\left\{n\geq 0:X^{n+L_{0}+L_{1}+\cdots+L_{m-1}}_{m}=0\right\},\qquad m\geq 1;$ (150)
also
$K_{0}=\inf\left\{n\geq 0:X^{n}_{n}=0\right\}$ (151)
and
$K_{m}=\inf\left\{n\geq 0:X^{n+K_{0}+K_{1}+\cdots+K_{m-1}}_{n-m}=0\right\},\qquad m\geq 1.$ (152)
Observe that
$X^{n+L_{0}+L_{1}+\cdots+L_{m}}_{m}=X^{n+K_{0}+K_{1}+\cdots+K_{m}}_{n-m}=0\qquad\text{for all }n\geq 0.$ (153)
Each of the above random variables resembles the variable $L$ of the appendix. The boundary condition $X^{0}_{0}=1$ implies that $\mathbb{P}\{L_{0}=0\}=\mathbb{P}\{K_{0}=0\}=0$ and, hence, formula (172) of the appendix together with (4) imply that
[…]
and
[…]
Regarding $L_{m}$ and $K_{m}$, $m\geq 1$, formulas (4), (171), and (172) imply that
$\mathbb{P}\left\{L_{m}=0\right\}=\mathbb{P}\left\{X^{L_{0}+L_{1}+\cdots+L_{m-1}}_{m}=0\right\}$ (156)
and
$\mathbb{P}\left\{L_{m}=k\right\}=\mathbb{P}\left\{X^{L_{0}+L_{1}+\cdots+L_{m-1}}_{m}=1\right\}(1-\beta)\beta^{k-1},\qquad k\geq 1,$ (157)
while (symmetrically)
$\mathbb{P}\left\{K_{m}=0\right\}=\mathbb{P}\left\{X^{K_{0}+K_{1}+\cdots+K_{m-1}}_{n-m}=0\right\}$ (158)
and
$\mathbb{P}\left\{K_{m}=k\right\}=\mathbb{P}\left\{X^{K_{0}+K_{1}+\cdots+K_{m-1}}_{n-m}=1\right\}(1-\alpha)\alpha^{k-1},\qquad k\geq 1.$ (159)
Notice that the random variables $L_{m}$, $m\geq 0$, as well as the variables $K_{m}$, $m\geq 0$, are not independent.
Example 5. Let us calculate $\mathbb{P}\left\{L_{1}=0\right\}$. In view of (149), (154), and (156) (as well as (4) and the independence of $X_{0}^{1}$ and $X^{1}_{1}$) we have
$\mathbb{P}\left\{L_{1}=0\right\}=\mathbb{P}\left\{X^{L_{0}}_{1}=0\right\}=\sum_{k=1}^{\infty}\mathbb{P}\left\{X^{L_{0}}_{1}=0\,|\,L_{0}=k\right\}\mathbb{P}\left\{L_{0}=k\right\}$
$\quad=\mathbb{P}\left\{X^{1}_{1}=0\,|\,X_{0}^{1}=0\right\}(1-\beta)+(1-\beta)\sum_{k=2}^{\infty}\mathbb{P}\left\{X^{k}_{1}=0\,|\,L_{0}=k\right\}\beta^{k-1}$
$\quad=(1-\alpha)(1-\beta)+(1-\beta)\sum_{k=2}^{\infty}\mathbb{P}\left\{X^{k}_{1}=0\,|\,L_{0}=k\right\}\beta^{k-1}.$ (160)
Now, in view of (150) we have that
$\mathbb{P}\left\{X^{k}_{1}=0\,|\,L_{0}=k\right\}=\mathbb{P}\left\{X^{k}_{1}=0\,|\,X_{0}^{1}=X_{0}^{2}=\cdots=X_{0}^{k-1}=1,X^{k}_{0}=0\right\}.$ (161)
Proceeding as in Example 4 we get
$\mathbb{P}\left\{X^{k}_{1}=0\,|\,X_{0}^{1}=X_{0}^{2}=\cdots=X_{0}^{k-1}=1,X^{k}_{0}=0\right\}$
$\quad=(1-\alpha)w_{k-1}+(1-\gamma)(1-w_{k-1})=(\gamma-\alpha)w_{k-1}+1-\gamma,$ (162)
where $w_{k}$ is given by (29).
Finally, by using (162), (161), and (29) in (160) we obtain (by summing the resulting geometric series)
$\mathbb{P}\left\{L_{1}=0\right\}=\frac{1-(\gamma-\alpha)\beta-\alpha}{1-(\gamma-\alpha)\beta}.$ (163)
$\diamondsuit$
6 Epilogue
The model presented in this paper can be extended/enriched in various ways. For instance:
(i) We can assume that the fire starts at many points, say at the points (j,1)(j,-1), $k\in F_{x}$ and (1,k)(-1,k), $k\in F_{y}$, where $F_{x},F_{y}\subset\mathbb{N}$. In this case, the “boundary conditions” for $X(j,k)$ are
$X(j,-1)=X(-1,k)=1,\quad j\in F_{x},\ k\in F_{y},$
and
$X(j,-1)=X(-1,k)=0,\quad j\in\mathbb{N}\smallsetminus F_{x},\ k\in\mathbb{N}\smallsetminus F_{y}.$
With such boundary conditions we can furthermore assume that X(1,1)=0X(-1,-1)=0 and, also, $X(j,k)=0$ whenever $j\leq-2$ or $k\leq-2$.
(ii) The status $X(j,k)$ of the tree at the point $(j,k)$ can take more than two values, say $X(j,k)\in\{0,1,\ldots,m\}$ to take into account the case where the tree at $(j,k)$ is partially burnt (i.e. burnt to a certain degree).
(iii) We can consider a higher dimensional model. For example we can consider a corrosion model: Each point $(j,k,\ell)\in\mathbb{N}^{3}$ is associated to a molecule (or atom, or ion) and the status $X(j,k,\ell)$ of the molecule at $(j,k,\ell)$ is affected only by the status of the neighboring molecules at $(j-1,k,\ell)$, $(j,k-1,\ell)$, and $(j,k,\ell-1)$.
(iv) Finally, one might investigate whether there is a continuous analog of the presented model.
7 APPENDIX: An one-dimensional model
Let us discuss briefly the (almost trivial) one-dimensional case, where the forest is represented by the set $\mathbb{N}=\{0,1,2,\ldots\}$ and each point $j\in\mathbb{N}$ is associated to a tree. The status of the tree at $j$ is $X(j)$, so that if $X(j)=1$, the tree at $j$ is burnt, while if $X(j)=0$, then the tree at $j$ is not burnt.
Here the model goes as follows: The status of the tree at $j$ is affected only by the status of the neighboring tree at j1j-1 so that
$\mathbb{P}\{X(j)=1\,|\,X(j-1)=1\}=p\qquad\text{and}\qquad\mathbb{P}\{X(j)=1\,|\,X(j-1)=0\}=0,$ (164)
where
$0\leq p\leq 1.$
The parameter (probability) $p$ depends on the strength of the wind. The smaller the magnitude of the velocity of the wind, the closer to 0 is pp.
We assume that the fire starts at the point 0 with a given probability $r\in[0,1]$. Thus the boundary condition for $X(j)$ is
$\mathbb{P}\{X(0)=1\}=r$ (165)
(this is slightly more general than the condition $X(0)=1$).
Notice that our assumptions imply
$\{X(j)=1\}\subset\{X(j-1)=1\},$ (166)
hence
$\{X(j)=1\}=\{X(0)=1\}\cap\{X(1)=1\}\cap\cdots\cap\{X(j)=1\}.$ (167)
From (164) and (165) we get that
$\mathbb{P}\{X(1)=1\}=\mathbb{P}\{X(1)=1\,|\,X(0)=1\}\mathbb{P}\{X(0)=1\}=rp,$
$\mathbb{P}\{X(2)=1\}=\mathbb{P}\{X(2)=1\,|\,X(1)=1\}\mathbb{P}\{X(1)=1\}=rp^{2},$
and, in general,
$\mathbb{P}\{X(j)=1\}=rp^{j},\qquad j\geq 0.$ (168)
Let
$L=\inf\left\{n\geq 0:X(n)=0\right\},$ (169)
so that
$X(n+L)=0\qquad\text{for all }n\geq 0.$ (170)
In fact, $L$ is the (total) number of burnt trees.
From (165) and (168) we get
$\mathbb{P}\{L=0\}=\mathbb{P}\{X(0)=0\}=1-r$ (171)
and
$\mathbb{P}\{L=k\}$ $=\mathbb{P}\{X(k)=0,X(k-1)=1\}$
$=\mathbb{P}\{X(k)=0\,|\,X(k-1)=1\}\mathbb{P}\{X(k-1)=1\}$
$=r(1-p)p^{k-1},\qquad k\geq 1.$ (172)
Notice that the conditional distribution of the random variable $L$ given $\{X(0)=1\}$ is geometric with parameter 1p1-p. Thus, unless $p=1$ in which case the whole forest will be burnt, we have that $L<\infty$ almost surely (which means that most of the forest will be spared).
Finally, by a direct calculation, or by conditioning on $\{X(0)=0\}$ and $\{X(0)=1\}$ and by invoking the properties of the geometric distribution we easily obtain that
$\mathbb{E}[L]=\frac{r}{1-p}\qquad\text{and}\qquad\mathbb{V}[L]=\frac{r(1+p-r)}{(1-p)^{2}}.$ (173)